Michael is two years old than Kyle. Twenty-two years ago, kyles age was four-fifths of michaels. How old is each now?
M=k+2
4/5(m-22)=k-22

Answer by jorel1380(3719)

m=k+2
4/5(m-22)=k-22
4/5((k+2)-22)=k-22
4/5(k-20)=k-22
4k-80=5k-110
k=30
m=32
☺☺☺☺